OpenStudy (anonymous):

f={(0,5),(1,4),(2,7),(3,14)} g(x)=3f(-x+1)+7 a) evaluate g(-1) somehow the textbook got an answer of 28 b) evaluate g(1) textbook had an answer of 22 c) evaluate g(3) textbook had an answer of undefined

OpenStudy (anonymous):

for each of these, note what you are substituting for x in g(x), and substitute the same for f(x): a) g(-1) = 3f(-(-1) + 1) + 7 = 3f(1+1) + 7 = 3f(2) + 7 = 3(7) + 7 = 28 b) g(1) = 3f(-(1) + 1) + 7 = 3f(0) + 7 = 3(5) + 7 = 22 c) g(3) = 3f(-(3) + 1) + 7 = 3f(-2) + 7. The answer is undefined because f(-2) is undefined.

OpenStudy (anonymous):

oh now i get it =) thank you

OpenStudy (anonymous):

ive been trying to figure this out for an hour, thank you so much

OpenStudy (anonymous):

no problem